Circle one pair of like terms and underline the other (2)


a) 4r – 6 = -13r + 11 b) 2.5 q + 7 = 1.6 q


1
Expert's answer
2021-04-20T02:11:34-0400

a) 4r - 6 = -13r + 11

4r + 13r = 11 + 6

4r + 13r (like terms)

"4r +13r = \underline{11+ 6}"


b) 2.5q + 7 = 1.6q

2.5q - 1.6q = -7

2.5q - 1.6q (like terms)

"2. 5q - 1.6q = \underline{-7}"